9 of 9 Problema #6 The breakdown of executed instructions is as follows: add addi 30% not 5% Iw 115% 10% 135% 15% 1. In what fraction of all cycles is the data memory used? 2. In what fraction of all cycles is the output of the sign-extend circuit needed?
Answer #1

1. Data memory is used only in Load and Store instruction, which are "lw" and "st" instruction here. So out of 100 % instruction, "lw" and"st" instruction are 15+5 = 20%

So fraction of cycle used for data memory = 20/100 = 1/5

2. Sign extended circuit is needed only for addi, lw and sw instruction because these are the only operation in which some immediate value is present in the instruction and to add a value with immediate operand, sign extension is needed.

So total percentage of instruction which need sign extension are = 30 + 15 + 5 = 50 %

So fraction of cycle used for data memory = 50/100 = 1/2

Please comment for any clarification.